OpenStudy (anonymous):

Can anyone help with finding the following limit? It's in my FP3 textbook and the answer is said to be 3/2, but no method is shown. The method required involves Maclaurin's theorem, and I assume the standard expansion of (1+x)^n is used.

OpenStudy (anonymous):

\[\lim_{x \rightarrow \infty}\left[ \left( x^2 + 3x \right)^\frac{ 1 }{ 2 } - x\right]\]